OpenStudy (anonymous):

Find the percent increase or decrease from 60 to 600.

OpenStudy (anonymous):

percent change is (new amount - old amount) / old amount. So, here, your new amount is 600 and your old amount was 60.

OpenStudy (anonymous):

So that would be: (600 - 60)/60 Make sure you convert that decimal to a percentage...

OpenStudy (anonymous):

900%?

OpenStudy (anonymous):

A) 850% B) 900% C) 100% D) 950%

OpenStudy (anonymous):

those are m answer choices

OpenStudy (anonymous):

yes :) I didn't mean to confuse it by saying "convert the decimal"... didn't think about the fact that it would be an even 9. 900% is correct

OpenStudy (anonymous):

coud you help me on this one too? x.x

OpenStudy (anonymous):

Find the percent increase or decrease from 65 to 120.

OpenStudy (anonymous):

It's the same idea: (new - old) / old = (120 - 65)/65

OpenStudy (anonymous):

0.846

OpenStudy (anonymous):

Is the answer to that but do i round it?

OpenStudy (anonymous):

87.5% 65% 84.30% 84.62% are my answer choices

OpenStudy (anonymous):

you need to have it as a percentage... 0.84615 = 84.615 % But yes, then you round it appropriately based on your answers.

OpenStudy (anonymous):

So, 84.62%

OpenStudy (anonymous):

yes... correct :)

OpenStudy (anonymous):

Thanks! Could you help me with a few more? :)

OpenStudy (anonymous):

Sure, glad to help...

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

The cost of sugar rises from $20 per kg to $25 per kg. What is the percentage increase for the cost of sugar? 25% 50% 30% 20%

OpenStudy (anonymous):

can you set it up this time? Percent increase = (new - old)/old =

OpenStudy (anonymous):

(20-25)/20 ??

OpenStudy (anonymous):

like that?

OpenStudy (anonymous):

Close... you probably want to subtract 20 from 25... like 25 - 20 rather than 20-25

OpenStudy (anonymous):

(current price - previous price) / previous price = (25 - 20) / 20

OpenStudy (anonymous):

25%?

OpenStudy (anonymous):

correct :)

OpenStudy (anonymous):

if you had done 20-25 / 20, you would have gotten -25%... you could just realize that you subtracted in the wrong direction... that an increase from 20 to 25 should be a positive percentage.

OpenStudy (anonymous):

Oh,

OpenStudy (anonymous):

The biggest cause of error on these is accidentally dividing by the new amount instead of the old amount. you find percentage change by the amount of change divided by the old amount.

OpenStudy (anonymous):

but you didn't make that error :)

OpenStudy (anonymous):

x) yay

OpenStudy (anonymous):

The cost of a car is $35,000. If the cost is increased by $7000, what is the percent increase? 20% 25% 50% 2%

OpenStudy (anonymous):

yay... small victories are nice :)

OpenStudy (anonymous):

go for it... I'll check your work.

OpenStudy (anonymous):

this one's slightly different, but it's the same idea... how much did it change, and then divide the change amount by the original amount.

OpenStudy (anonymous):

(35000-7000)/7000

OpenStudy (anonymous):

no, that's why this one is a little different. the price is 35000 They don't tell you the new price... they just say it went up by 7000. So you could add and find new price as 35000 + 7000 = 42000 Then (new - old) / old = (42000 - 35000) / 7000 But you can also just realize that the stuff in the numerator is the "amount of change"... and here, they just told you that the amount of change is 7000. So you can find percentage change by (amount of change) / (original amount) 7000/35000

OpenStudy (anonymous):

5000

OpenStudy (anonymous):

Oh wait 0.2 or 2%

OpenStudy (anonymous):

sorry, I made a typo in the middle up there... it should have been Then (new - old) / old = (42000 - 35000) / 35000 oops!

OpenStudy (anonymous):

7000/35000 = 1/5 = 20% That's right, glad you caught it :)

OpenStudy (anonymous):

The cost of a dress is $ 150. If its cost is increased by 10%, what is the new cost? (I'm timed and didn't realize it. I have 20 minutes left) D:

OpenStudy (anonymous):

Same idea: multiply 150 * 10%... that's the increase amount. Then add to $150 to get the new cost.

OpenStudy (anonymous):

2400?

OpenStudy (anonymous):

choices: A) $ 155 B) $ 165 C) $ 166 D) $ 160

OpenStudy (anonymous):

150 * 10% = 150 * 0.10 = $15 increase Old cost: $150 New cost: $150 + increase = $150 + $15

OpenStudy (anonymous):

Got it! And 3 more questions okay?

OpenStudy (anonymous):

Tom's commute was 90 miles, but it increased by 40% when he changed jobs. What is the distance of his new commute?

OpenStudy (anonymous):

Same exact thing: Increase amount = 40% * 90 miles = 36 miles New commute = 90 + 36

OpenStudy (anonymous):

126 miles

OpenStudy (anonymous):

(10 mins left@ :( ) The cost of a dress is $120. If its cost is increased by 5%, what is the new cost?

OpenStudy (anonymous):

Same again: increase amount = 120 * 0.05 = 6 Old amount: 120 New amount 120 + 6 You can get these I think... :) Too bad they rush you.

OpenStudy (anonymous):

Ikr. I hate being rushed. -.- I can't learn properly. :(

OpenStudy (anonymous):

The cost of a wallet is $120. If it's cost decreases by 15%, then calculate the new cost.

OpenStudy (anonymous):

cost decrease: 15% * 120 = 18 Old cost: 120 New cost: 120 - 18

OpenStudy (anonymous):

5 mins. x.x What speed is a car traveling if there is a 25% decrease in the speed of a car whose original speed was 80 miles/hr?

OpenStudy (anonymous):

decrease: 80 * .25 = 20 old: 80 new: 80 - 20

OpenStudy (anonymous):

Find the percent increase or decrease from 25 to 100.

OpenStudy (anonymous):

last one!

OpenStudy (anonymous):

I know its not 75% because theres no answer choice for that. XD I just can't get this last one. x.x

OpenStudy (anonymous):

should be 400% increase, right?

OpenStudy (anonymous):

i think. not sure. x.x the answer choices are A) 350% B) 625 C) 300% D) 325%

OpenStudy (anonymous):

4 times as much?

OpenStudy (anonymous):

no, 100-25/25 = 75 / 25 = 300%

OpenStudy (anonymous):

sorry, this pace is a recipe for error :(

OpenStudy (anonymous):

tanks

OpenStudy (anonymous):

done?